# Quantum transistors for heat flux in and out of working substance parts: harmonic vs transmon and Kerr environs

### v1

@@ II.3 Relevant heat currents in working-substance thermal transistors
where $X\in\{L,M,R\}$. The operator, $\dot{\rho}_{X}$, denotes the rate of change of the reduced density matrix of the $X^{\text{th}}$ terminal of the working substance. This type of heat current can be considered “local” in a sense that heat flows locally in and out of a particular subsystem of the working substance. The difference between the two approaches is pictorially depicted in Fig 1.
@@
Figure 2: (a) Variation of $\partial J_{X}/\partial T_{M}$ with $T_{M}$, for $X\in\{L,M,R\}$, is depicted. Values of $\partial J_{M}/\partial T_{M}$ are much smaller than the other two, and becomes equal to zero at a particular value of $T_{M}$ which is equal to $6.65$ in units of $\tilde{T}$ when $g=4$. (b) Variation of $\partial{J_{M}}/\partial{T_{M}}$ with change in the coupling constant between the qubit and the environment is demonstrated. We see that, for every value of $g$, the quantity $\partial{J_{M}}/\partial{T_{M}}$ becomes zero at some value of $T_{M}=T_{M}^{critical}$, which makes the amplification factor undefined. This value of $T_{M}^{critical}$ increases as we increases the value of $g$. The black dotted line is for $\partial{J_{M}}/\partial{T_{M}}=0.$ The horizontal axis has dimension $\tilde{T}$ while the vertical axis when multiplied by $\hbar/\tilde{T}$ will have the dimension of ratio of rate of change of energy with temperature.
@@ III WTT in presence of linear environment
In this section, we work with linear environment and consider the parameter space which allows the WIT to produce amplification. The temperature of the middle environment is taken to be in between that of the left and right environments. We consider the initial state of the system comprising the three qubits to be $|000\rangle\langle 000|$, where |$\ket{0}$ denotes the eigenstate of Pauli-$z$ with eigenvalue $+1$. The initial state of the environment $X$ is considered to be the thermal state, given by
@@
We have defined time as $t\leavevmode\nobreak\ \tilde{t}$ where $t$ is a dimensionless quantity whose magnitude denotes the magnitude of time, and $\tilde{t}$ has the dimension of time with magnitude equal to $1$. Similarly, temperature is defined as $T\leavevmode\nobreak\ \tilde{T}$ where $T$ is dimensionless quantity denoting the magnitude of temperature, and $\tilde{T}$ has the dimension of temperature with magnitude equal to $1$. $\tilde{t}$ and $\tilde{T}$ are related by the relation
$\tilde{T}=\frac{\hbar}{k_{B}\leavevmode\nobreak\ \tilde{t}},$
@@ III.2.1 WTTs with three qubits and two environments
III.2.2 Amplification with two qubits and two environments
where $J_{L}$ and $J_{R}$ are the heat currents flowing into the left and right qubits respectively. In this case, we observe that when the energy difference between the ground state and first excited state is the same for left and right qubit, then there is no amplification. However, when we consider unequal energy differences, there is a finite amplification in a small range of $T_{L}$. The range of $T_{L}$ is from 0 to 2.488 when $\omega_{L}=1$ and $\omega_{R}=2$, $T_{R}=4,g=4,\Delta=5$ and $t=1$.
@@ IV Response to non-linear environments in amplification of WTT
We further consider how the variation of amplification with time gets affected when non-linearity is introduced in the system. For the transmon and Kerr type environments, we observe that the amplification as a function of time, follows a trend which is similar to that of the linear environment case. However, the change in amplification when we consider non-linear environments is very small ($\sim 10^{-2}$) at all points of time except at time $t=\tilde{t}$, where the amplification has significant differences, i.e. $\epsilon=-0.01$ gives an amplification of $38.98$, $\epsilon=0.01$ gives an amplification of $34.00$, meanwhile linear type of environment gives an amplification of $36.27$. The maximum amplification that one can attain using this setup in presence of Kerr type environments is slightly higher $(38.98)$ than that using transmon qutrits $(34.00)$, albeit an amplification close to the linear case is obtained in both the cases.
@@ V Quantifying non-Markovianity
As previously mentioned, the system that we consider deals with the global unitary evolution of the system and environment, followed by discarding the environment. This is, in general, non-Markovian in nature since there are no assumptions of Markovianity. Markovinity is a special case that arises when the environment does not possess the memory of its past events, i.e. due to a memoryless environment. In our case, when the environment interacts with the system, the memory of the environment changes depending upon the interaction. So this dynamics is non-Markovian at this stage, since the environment evolves depending upon its previous memory. However, the environment returns to its initial state periodically, i.e. there is also a loss of memory of the environment, which may potentially lead to Markovian behavior. Hence there is a tussle between memory-dependence and memory-independence of the environment in the transistor model that we have chosen. In this section, we study the memory-effects of the environment on the system, and quantify it in terms of the non-Markovinity induced in the system.
Different measures have been defined to detect non-Markovinity of a system undergoing a certain dynamics. We use the Breuer-Laine-Piilo (BLP) measure Breuer et al. (2009) - a faithful measure of non-Markovinity - to quantify non-Markovinity in our system. The trace distance, $D(\rho_{1},\rho_{2})$, between two states, $\rho_{1}$ and $\rho_{2}$, is a measure of how distinguishable the two states are from each other. It is defined by
@@
To draw a parallel with the trends in amplification of the heat currents associated with each qubit, we consider the measure of non-Markovinity for each individual qubit of the working substance. We plot the measure $\mathcal{N}(\Phi)$ for each qubit, along the vertical axis, versus time $t$, along the horizontal axis in Fig. 8. In order to do so, we optimize over the state-space comprising of single qubits. It is to note that the optimal state-pairs belong to the boundary of the state-space Wißmann et al. (2012). Further, since we work with qubits, the boundary of the state-space contains only pure states, and therefore it is sufficient to optimize over pure states only. We consider pure states of the form
$\rho=\frac{1}{2}(\mathbb{I}+\vec{a}\leavevmode\nobreak\ \cdot\leavevmode\nobreak\ \vec{\sigma}),$ (10)
@@
Figure 8: Plot of the BLP measure, $\mathcal{N}(\Phi)$, for each qubit of the working substance versus time, $t$. Here the dark blue, light blue, and grey curves correspond to the amplifications for the left, middle and right qubits of working substance respectively. We find that the non-Markovian effect exists up to time $t=1.5\tilde{t}$, after which the system behaves as Markovian.
@@
Noting that the system is given periodic kicks of Markovianity through successive collisions between the system and environment, it is difficult to ascertain apriori, whether non-Markovian dynamics will dominate over Markovian. In Fig. 8, we observe the BLP measure for each qubit of the working substance, as a function of time. We find that at shorter timescales, there is back flow of information from the environment, and system is non-Markovian, as indicated by non-zero value of $\mathcal{N}(\Phi)$. However, at longer timescales, viz. after time $1.5\tilde{t}$, the BLP measure $\mathcal{N}(\Phi)=0$, implying that the system has reached a Markovian stage. This happens as the system loses memory of non-Markovian interactions at larger times. However, this Markovian feature is not reflected in the amplification (refer to Fig. 7), since there are non-monotonic jumps in the amplification even after time $1.5\tilde{t}$. This is because certain quantities, like the amplification in our case, has greater memory than the measure of non-Markovinity that we use here. The trend in the behavior of $\mathcal{N}(\Phi)$ is qualitatively the same for the left, middle and right qubits, as depicted in Fig. 8.
VI Conclusion

### v2

@@ II.3 Relevant heat currents in working-substance thermal transistors
where $X\in\{L,M,R\}$. The operator, $\dot{\rho}_{X}$, denotes the rate of change of the reduced density matrix of the $X^{\text{th}}$ terminal of the working substance. This type of heat current can be considered “local” in a sense that heat flows locally in and out of a particular subsystem of the working substance. The difference between the two approaches is pictorially depicted in Fig 1. This definition of heat current is constructed in a generalized setting, and is not derived via any master equation formalism.
In the fabrication of quantum devices, the material properties of the qubits, such as melting point, are typically well-characterized in advance and require minimal information about the surrounding environment. From this standpoint, analyzing heat currents from the perspective of the qubits themselves, rather than from the conventional bath-centric viewpoint, enables a direct correlation between these currents and the intrinsic material properties that are readily accessible in experiments. Such a system-focused definition of heat current has the potential to guide the design and optimization of qubit-based quantum devices whose performance is sensitive to the behavior of these currents.
In our work, the dynamics of the working substance is modeled as general CPTP dynamics, realized through global unitary evolution of the combined system and environment, followed by tracing out the environment. This framework goes beyond the Markovian assumption. Therefore our approach does not rely on master equations, neither global nor local and as the system does not relax to a steady state. Our numerical analysis further confirms that the system exhibits continuous time-dependent evolution without approaching a steady state. Consequently, since no steady state is attained, the currents do not sum to zero as in their classical counterparts. Within this setting, the definition of heat current that we introduce is novel, as it captures the flow of heat through the system even in the absence of a steady state.
@@
Figure 2: (a) Variation of heat currents at time = 1 $\tilde{t}$ (b) Variation of $\partial J_{X}/\partial T_{M}$ with $T_{M}$, for $X\in\{L,M,R\}$, is depicted. Values of $\partial J_{M}/\partial T_{M}$ are much smaller than the other two, and becomes equal to zero at a particular value of $T_{M}$ which is equal to $6.65$ in units of $\tilde{T}$ when $g=4$. (c) Variation of $\partial{J_{M}}/\partial{T_{M}}$ with change in the coupling constant between the qubit and the environment is demonstrated. We see that, for every value of $g$, the quantity $\partial{J_{M}}/\partial{T_{M}}$ becomes zero at some value of $T_{M}=T_{M}^{critical}$, which makes the amplification factor undefined. This value of $T_{M}^{critical}$ increases as we increases the value of $g$. The black dotted line is for $\partial{J_{M}}/\partial{T_{M}}=0.$ The horizontal axis has dimension $\tilde{T}$ while the vertical axis when multiplied by $\hbar/\tilde{T}$ will have the dimension of ratio of rate of change of energy with temperature.
@@ III WTT in presence of linear environment
[…]
@@
We have defined time as $t\penalty 10000\ \tilde{t}$ where $t$ is a dimensionless quantity whose magnitude denotes the magnitude of time, and $\tilde{t}$ has the dimension of time with magnitude equal to $1$. Similarly, temperature is defined as $T\penalty 10000\ \tilde{T}$ where $T$ is dimensionless quantity denoting the magnitude of temperature, and $\tilde{T}$ has the dimension of temperature with magnitude equal to $1$. $\tilde{t}$ and $\tilde{T}$ are related by the relation
$\tilde{T}=\frac{\hbar}{k_{B}\penalty 10000\ \tilde{t}}$,
@@ III.1.1 Dependence of amplification on the system-environment interaction
In Fig. 2(a), we have now added a curve where the heat current flowing through the $X^{\text{th}}$ qubit, $J_{X}$, is plotted along the vertical axis, as a function of the middle-bath temperature, $T_{M}$ along the horizontal axis. Here we see that compared to the variation in $J_{L}$ and $J_{R}$, the quantity $J_{M}$ varies very slowly with respect to $T_{M}$. This feature is also evident from Fig. 2(b), where see that the quantity, $\frac{\partial J_{M}}{\partial T_{M}}$}, is much smaller than $\frac{\partial J_{L}}{\partial T_{M}}$} and $\frac{\partial J_{R}}{\partial T_{M}}$}. Therefore, the denominator of the dynamical amplification factor becomes significantly higher than its numerator, resulting in an appreciable amplification in the parameter regime in which we work. Let us now qualitatively interpret the behavior of the amplification factor, $\alpha_{L}$, as given in Fig. 3, by looking at the heat-current derivatives given in Figs. 2(b) and 2(c). From Fig. 2(c), we see that in the temperature regime, $T_{M}<T^{critical}_{M}$, the quantity, ${\frac{\partial J_{M}}{\partial T_{M}}}<0$. Then it gradually increases to $T_{M}=T^{critical}_{M}$, where ${\frac{\partial J_{M}}{\partial T_{M}}}=0$ and in the regime $T_{M}>T^{critical}_{M}$, the quantity, ${\frac{\partial J_{M}}{\partial T_{M}}}>0$. Further, from Fig. 2(b) we observe that the quantity, $\frac{\partial J_{L}}{\partial T_{M}}>0$, $\forall$ $T_{M}$. Therefore, for small values of $T_{M}$, the value of $\frac{\partial J_{L}}{\partial T_{M}}$} is positive and $\frac{\partial J_{M}}{\partial T_{M}}$} is negative, which causes negative amplification. At $T_{M}=T^{critical}_{M}$, the quantity, ${\frac{\partial J_{M}}{\partial T_{M}}}=0$, which causes the amplification factor, $\alpha_{L}$, to diverge at the critical value of $T_{M}=T^{critical}_{M}$. Finally, for values of temperatures of the middle bath, $T_{M}>T^{critical}_{M}$, both $\frac{\partial J_{M}}{\partial T_{M}}$} and $\frac{\partial J_{L}}{\partial T_{M}}$} are positive, which results in positive amplification.
@@ III.1.2 Variation of amplification with time
In order to understand why there is a peak in the amplification versus time plot at $t=1\tilde{t}$, we carried out a perturbative analysis of how the amplification factor behaves around time $t=1$ within the relevant parameter regime. This revealed some interesting features specific to the Hamiltonian considered in our work. In particular, by varying the coupling strength between each qubit and its respective bath, we observed the following. For coupling strengths slightly below $g=4$, the amplification is small and positive, whereas for coupling strengths slightly above $g=4$, it becomes small and negative. The magnitude of the amplification exhibits a sharp peak at $g=4$, where we observe a pronounced amplification, as illustrated in Fig. 5. This behavior is an intrinsic feature of the Hamiltonian model in the chosen parameter regime.
Figure 5: Variation of $\alpha_{L}$ on the vertical axis as the coupling strength $g$ (on the horizontal axis) between the qubit and the environment is varied at $t=1\tilde{t}$. We observe that when the coupling strength $g<4$, the amplification is small and positive, and when $g>4$, it is small and negative. A sharp peak in amplification occurs around $g=4$, indicating a significant enhancement at this point.
@@ III.2.1 WTTs with three qubits and two environments
Figure 7: Variation of dynamical amplification factor with coupling strength $g$ when one of the environments is removed. The solid lines are for the system when the right environment is removed while the dotted lines are for the system when the left environment is removed .
@@
The height of the peak of $\alpha_{R}$ and $\alpha_{L}$ in Fig. 6 at time $t=0.7\tilde{t}$ can be explained by analyzing the variation of $\alpha$ with the coupling strength $g$ as shown in Fig. 7. We see that in the case where right environment is removed (corresponding to solid pink and orange lines), the amplification is very small and monotonously increases until it reaches its peak at $g=4.05$ and then suddenly becomes negative. After which the amplitude of the amplification keeps on decreasing. A similar trend is seen when the left environment is removed (corresponding to the dotted black and blue lines). In this case, the peak is reached at a slightly higher $g$ value of $g=4.2$ but the trend remains similar to that of the case when right environment was removed.
III.3 Symmetric versus asymmetric coupling between working-substance qubits
In the preceding subsections, we considered the case where there was no coupling between the left and right qubits of the working subsystem, and the couplings between the left–middle and right–middle qubits were equal, i.e
$\omega_{ML}$ $=\omega_{MR}=\Delta$
$\omega_{LR}$ $=0.$ (8)
Here we consider the following two situations. In the first case, we additionally couple the left and right qubits, and consider the coupling between the left–middle, right–middle and left–right qubits to be equal, i.e.
$\omega_{ML}=\omega_{MR}=\omega_{LR}=\Delta.$ (9)
In the second case, we consider unequal couplings between each pair of qubits, i.e,
$\omega_{ML}$ $=\Delta$
$\omega_{MR}$ $=\Delta+\delta$
$\omega_{LR}$ $=\Delta-\delta,$ (10)
where $\delta$ is a parameter considered small compared to $\Delta$. We refer to this case as the asymmetric case. In particular, we have taken $\delta=0.1$ and $\Delta=3$ as before.
We calculate the dynamical amplification factor for the left and right qubits as functions of the middle bath temperature and the evolution time, for both symmetric and asymmetric coupling scenarios, and present the results in Fig. 8. Our analysis shows that amplification persists under both types of coupling. The dependence of amplification on the middle bath temperature $T_{M}$, shown in Figs. 8 (a) and (c) corresponding to symmetric and asymmetric cases, follows a trend similar to that observed previously for the coupling strengths given in Eq. III.3 (refer to Fig. 3). Specifically, the amplification is initially negative, with its magnitude increasing up to a critical temperature $T_{M,A(S)}^{\text{critical}}$, beyond which it becomes positive and subsequently decreases. The value of $T_{M}^{\text{critical}}$ differs between the symmetric and asymmetric cases, with the critical temperature for symmetric case being $T_{M,S}^{\text{critical}}=1.75$ and for the asymmetric case being $T_{M,A}^{\text{critical}}=10.45$
The time dependence of the amplification, shown in Figs. 8 (b) and (d), also exhibits distinct features. After $t=\tilde{t}$, the amplification becomes periodic with the same period as the interaction time, consistent with the behavior reported in Fig. 4. In both symmetric and asymmetric cases, the amplification reaches its peak at $t=0.4\tilde{t}$.
Figure 8: Variation of amplification $\alpha_{X}$ as temperature of the middle bath $T_{M}$ and time $t$ change. The top figures (a) and (b) correspond to the symmetric case where the coupling between all the qubits is $\Delta$. The bottom figures correspond to the asymmetric coupling between the pair of qubits. The coupling between left and middle is $\Delta=3$, between middle and right is $\Delta+0.1=3.1$ and between left and right is $\Delta-0.1=2.9$. (a) and (c) corresponds to time $0.4\tilde{t}$ while (b) and (d) correspond to $T_{M}=10$
@@ IV Response to non-linear environments in amplification of WTT
We further consider how the variation of amplification with time gets affected when non-linearity is introduced in the system, and demonstrate the results in Fig. 10. For the transmon and Kerr type environments, we observe that the amplification as a function of time, follows a trend which is similar to that of the linear environment case. However, the change in amplification when we consider non-linear environments is very small ($\sim 10^{-2}$) at all points of time except at time $t=\tilde{t}$, where the amplification has significant differences, i.e. $\epsilon=-0.01$ gives an amplification of $38.98$, $\epsilon=0.01$ gives an amplification of $34.00$, meanwhile linear type of environment gives an amplification of $36.27$. The maximum amplification that one can attain using this setup in presence of Kerr type environments is slightly higher $(38.98)$ than that using transmon qutrits $(34.00)$, albeit an amplification close to the linear case is obtained in both the cases.
Figure 11: Variation of amplification ($\alpha_{L}$) versus the temperature of the middle bath ($T_{M}$). The numerical analysis is done at time $t=0.4\tilde{t}$. Panel (a) corresponds to symmetric coupling between the qubits of the working substance i.e. $\omega_{ML}=\omega_{MR}=\omega_{LR}=\Delta$. The inset in the figure shows zoomed in area with $3.2<T_{M}<4.2$. Panel (b) corresponds to asymmetric coupling between the working substance qubits i.e. $\omega_{ML}=\Delta,\omega_{MR}=\Delta+\delta,\omega_{LR}=\Delta-\delta$, where $\delta=0.1$.
Next we introduce symmetric and asymmetric coupling between the qubits of the working substance along with non-linearities in the individual bath qutrits. The local Hamiltonian of the environment connected to each qubit is given by Eqn. 11 as discussed in the previous section. Meanwhile, the coupling between working-substance qubits are given by Eqns. 9 and III.3 for symmetric and asymmetric coupling respectively. In particular, we have taken $\delta=0.1$. The numerical values of amplification $\alpha_{L}$ are plotted versus the temperature of the middle bath $T_{M}$ for the symmetric and asymmetric cases in Fig. 11 (a) and (b) respectively. We show that a critical temperature exists for both symmetric and asymmetric coupling, at which a discontinuity in amplification occurs. These critical points are denoted by $T^{\text{critical}}_{M,S}$ and $T^{\text{critical}}_{M,A}$ for symmetric and asymmetric coupling, respectively. The insets in Fig. 11 provides a magnified view of regions below and above the critical temperature. We find that in both such regions, transmon-type environment offers an enhanced amplification. This is particularly different in the previous set of coupling defined by Eq. III.3 where in transmon and Kerr both showed an advantage in different regions. However, in the asymmetric coupling case, the relative advantage of transmon- and Kerr-type environments follows the same qualitative behavior as observed earlier. Specifically, in the low-temperature regime (to the left of the discontinuity, i.e. $T_{M}<T^{critical}_{M,A}$), Kerr-type non-linearity yields stronger amplification compared to the linear case, while in the high-temperature regime (to the right of the discontinuity, i.e. $T_{M}<T^{critical}_{M,A}$), transmon qutrits provide superior amplification relative to the linear case.
@@ V Quantifying non-Markovianity
As previously mentioned, the system that we consider deals with the global unitary evolution of the system and environment, followed by discarding the environment. This is, in general, non-Markovian in nature since there are no assumptions of Markovianity. markovianity is a special case that arises when the environment does not possess the memory of its past events, i.e. due to a memoryless environment. In our case, when the environment interacts with the system, the memory of the environment changes depending upon the interaction. So this dynamics is non-Markovian at this stage, since the environment evolves depending upon its previous memory. However, the environment returns to its initial state periodically, i.e. there is also a loss of memory of the environment, which may potentially lead to Markovian behavior. Hence there is a tussle between memory-dependence and memory-independence of the environment in the transistor model that we have chosen. In this section, we study the memory-effects of the environment on the system, and quantify it in terms of the non-markovianity induced in the system.
Different measures have been defined to detect non-markovianity of a system undergoing a certain dynamics. We use the Breuer-Laine-Piilo (BLP) measure Breuer et al. (2009) - a faithful measure of non-markovianity - to quantify non-markovianity in our system. The trace distance, $D(\rho_{1},\rho_{2})$, between two states, $\rho_{1}$ and $\rho_{2}$, is a measure of how distinguishable the two states are from each other. It is defined by
@@
To draw a parallel with the trends in amplification of the heat currents associated with each qubit, we consider the measure of non-markovianity for each individual qubit of the working substance. We plot the measure $\mathcal{N}(\Phi)$ for each qubit as a function of time $t$ in Fig. 12, where the vertical axis represents $\mathcal{N}(\Phi)$ and the horizontal axis represents $t$. Three scenarios are considered. Panel (a) corresponds to the case where the left–middle and right–middle qubits are coupled with the interaction strength specified in Eqn. III.3. Panels (b) and (c) illustrate the symmetric and asymmetric cases, respectively, with parameters chosen as described in Eqn. 9, III.3 respectively. In order to do so, we optimize over the state-space comprising of single qubits. It is to note that the optimal state-pairs belong to the boundary of the state-space Wißmann et al. (2012). Further, since we work with qubits, the boundary of the state-space contains only pure states, and therefore it is sufficient to optimize over pure states only. We consider pure states of the form
$\rho=\frac{1}{2}(\mathbb{I}+\vec{a}\penalty 10000\ \cdot\penalty 10000\ \vec{\sigma}),$ (13)
@@
Noting that the system is given periodic kicks of Markovianity through successive collisions between the system and environment, it is difficult to ascertain apriori, whether non-Markovian dynamics will dominate over Markovian. In all the panels of Fig. 12, we find that at shorter timescales, there is back flow of information from the environment, and system is non-Markovian, as indicated by non-zero value of $\mathcal{N}(\Phi)$. However, at longer timescales, viz. after time $1.5\tilde{t}$, the BLP measure $\mathcal{N}(\Phi)=0$, implying that the system has reached a Markovian stage. This happens as the system loses memory of non-Markovian interactions at larger times. However, this Markovian feature is not reflected in the amplification (refer to Fig. 4) corresponding to the parameter set considered in Fig. 12 (a), since there are non-monotonic jumps in the amplification even after time $1.5\tilde{t}$. This is because certain quantities, like the amplification in our case, have greater memory than the measure of non-markovianity that we use here. The trend in the behavior of $\mathcal{N}(\Phi)$ is qualitatively the same for the left, middle and right qubits for all three panels, as depicted in Fig. 12 (a), Fig. 12 (b) and Fig. 12 (c).
Figure 12: Plot of the BLP measure, $\mathcal{N}(\Phi)$, for each qubit of the working substance versus time, $t$ for (a) case with $\omega_{ML}=\omega_{MR}=\Delta,\omega_{LR}=0$, (b) symmetric coupling case, and (c) asymmetrical coupling case. Here the black, blue, and orange curves correspond to the amplifications for the left, middle and right qubits of working substance respectively.
VI Environment with Qubits
In this section, we aim to explore what happens when the environment consists of qubits instead of qutrits. Our primary objective is to determine whether a transistor effect persists when the working substance interacts with this type of environment.
Here we provide an analysis of the transistor action in a similar WTT model but with the environmental qutrits being replaced by qubits. In this regard, the local Hamiltonian of a single qubit of the environment gets modified to
$H_{env}=-\hbar\penalty 10000\ \Delta\sum_{i\in\{L,M,R\}}\sigma_{z}^{(i)},$ (14)
where $\sigma_{z}^{(i)}$ is the Pauli-$z$ matrix for spin-1/2 particles. Similarly, the interaction Hamiltonian also gets modified to
$H_{qubit-env}=-\hbar\penalty 10000\ g\sum_{i\in\{L,M,R\}}\sigma_{x}^{(i),qubit}\penalty 10000\ \sigma_{x}^{(i),qubit},$ (15)
where $\sigma_{x}^{(i),qubit}$ is the Pauli-$x$ matrix for spin-$1/2$ particle, and $g$ is the coupling constant between the system-qubit and the environment-qubit. The parameters considered are: $T_{L}=4,\penalty 10000\ T_{R}=10,\penalty 10000\ g=4,\penalty 10000\ \Delta=3,\penalty 10000\ T=5$.
We have plotted the dynamic curves for the dynamical amplification factors, $\alpha_{L}$ and $\alpha_{R}$, as functions of time $t$ in Fig. 13 (a). The maximum values of amplification for the left and right qubits, $\alpha_{L}$ and $\alpha_{R}$, are observed at time $9.7\tilde{t}$, with the corresponding maximum values being 37.46 and 73.67. At the same time, we have examined how the amplification associated with the left and right qubits changes as the temperature of the middle bath is varied. This behavior is illustrated in Fig. 13 (b), where it is shown that the amplifications for both qubits increase monotonically with an increase in the temperature of the middle bath. Hence, a two-level system will also be sufficient constituent of a collisional model as environment to observe transistor action.
Figure 13: In panel (a) Variation of the dynamical amplification factor, $\alpha_{L}$ and $\alpha_{R}$, with time, $t$. The vertical axis is dimensionless and the horizontal axis is in units of $\tilde{t}$. The given graph is at $T_{M}=10$ in units of $\tilde{T}$. At the same time in panel (b) Variation of the dynamical amplification factor, $\alpha_{L}$, along the vertical axis, versus the temperature of the middle bath, $T_{M}$, along the horizontal axis. The quantity, $\alpha_{L}$, is dimensionless, while $T_{M}$ is in units of $\tilde{T}$. The graph is plotted at time $t=9.7$.
VII Conclusion